# The immersed boundary method: an accelerated SIMPLE approach for moving bodies

### v1

@@ 5.3 Three-dimensional experiments: an oscillating porous sphere
A notable distinction between the two configurations lies in their flow penetration characteristics. The 7 sub-sphere array, with its higher porosity, exhibits more pronounced flow penetration between sub-spheres, manifested in elongated vortical structures that extend through the array (Fig. 9). In contrast, the 14 sub-sphere configuration shows more compact vortical structures that primarily develop around the array’s outer boundary (Fig. 10), consistent with its lower porosity. This difference is particularly evident at $Re=300$, where the vortical structures in the 7 sub-sphere case show distinct columnar formations that are absent in the denser 14 sub-sphere arrangement. Additionally, both configurations demonstrate asymmetric vortex shedding patterns between upward and downward motions, with more intense vortex formation occurring during direction reversal at the trajectory extrema. This asymmetry becomes more pronounced with increasing Reynolds number, particularly visible in the higher resolution structures at $Re=300$.
@@
After testing our accelerated SIMPLE-IBM on multiple sub-sphere arrays, we demonstrate that our solver’s performance scales independently of the number of sub-spheres. Table 4 shows the BiCGStab iteration count needed to solve Eq. (19) with preconditioning, averaged over a single oscillation period, for different Reynolds numbers (Re=50, 100, 200, and 300) and grid resolutions. The results show that this period-averaged iteration count remains nearly constant, requiring 4-5 iterations for both 7 and 14 sub-sphere configurations. Furthermore, these period-averaged iteration counts match the results in Table 2 for a single oscillating sphere, demonstrating consistent performance across different configurations. These numerical results, combined with Theorem 4.1, demonstrate our method’s applicability and scalability across a wide range of applications.
@@ 6 Conclusion and future work
We present a novel, scalable direct-forcing IBM for efficient moving boundary simulations that builds upon the SIMPLE-IBM method [16]. To address the pressure-forces coupling, a block reduction technique is applied to solve for the pressure corrections first, taking the Laplacian as a preconditioner for the resulting primal Schur complement. We prove rigorously and show numerically that the Laplacian is spectrally equivalent to the primal Schur complement. This key observation enables an efficient and robust method, as the pressure’s Laplacian, being independent of the coordinates of the immersed body, needs to be factorized only once—using a fast direct method [19]—and can then be reused throughout the simulation.
@@
We prove theoretically and demonstrate numerically that the efficiency of our approach does not depend on the grid resolution or on the number of moving bodies, thereby ensuring excellent scalability in time and memory consumption. This leads to notably low memory requirements, enabling precise moving boundary simulations (that often require high-performance computing facilities) to run on standard workstations or even laptops. Consequently, our approach enhances the accessibility of advanced moving boundary simulations to the CFD community. In this work, we examined the applicability of the method in a one-way coupled FSI framework. The methodology was successfully verified by simulating flows generated by a transversely oscillating sphere, showing good agreement with previously published results for various Reynolds numbers and oscillation amplitudes. The verified approach was then extended to more challenging configurations involving multiple moving bodies, specifically analyzing porous spheres modelled as arrays multiple sub-spheres. The simulations revealed distinct flow characteristics associated with different array porosities, including variations in drag coefficients, phase relationships, and vortical structure evolution, while maintaining computational stability and accuracy throughout all investigated parameters.
Future research directions should extend the method to two-way coupled FSI problems, specifically for the simulation of undulatory swimming and particle-laden flows. The efficiency of our preconditioner is independent of the specific discrete delta functions used, allowing for further improvements in accuracy through delta functions with a wider support [6]. As we approximate the regularization block by a scalar matrix, a key challenge lies in identifying delta functions with sufficiently wide support, for which the regularization block can still be well approximated by a scalar matrix.
Finally, our method employs a direct solver [19] designed for homogeneous boundary conditions. While this limitation is not particularly restrictive for our current application (as we use it to solve the Poisson equation for pressure corrections, typically formulated with homogeneous boundary conditions), there is potential for further generalization by applying, for example, a geometric multigrid [37] V-cycle to the Poisson equation in each preconditioning step, rather than a direct solution. Subtle fine-tuning of the multigrid method might also yield linear time scaling. The above-described enhancements would further increase the versatility of our approach for a wider range of FSI problems and enable better spatial accuracy.

### v2

@@ 6.1.2 Model problem 2: multiple packed spheres (one-way coupling)
A notable distinction between the two configurations lies in their flow penetration characteristics. The 7 sub-sphere array, with its higher porosity, exhibits more pronounced flow penetration between sub-spheres, manifested in elongated vortical structures that extend through the array (Fig. 7). In contrast, the 14 sub-sphere configuration shows more compact vortical structures that primarily develop around the array’s outer boundary (Fig. 8), consistent with its lower porosity. This difference is particularly evident at $Re=300$, where the vortical structures in the 7 sub-sphere case show distinct columnar formations that are absent in the denser 14 sub-sphere arrangement. Additionally, both configurations demonstrate asymmetric vortex-shedding patterns between upward and downward motions, with more intense vortex formation occurring during direction reversal at the trajectory extrema. This asymmetry becomes more pronounced with increasing Reynolds number, particularly visible in the higher resolution structures at $Re=300$.
6.1.3 Model problem 3: sedimentation and buoyant rise of a spherical particle (two-way coupling)
To further validate the accuracy and robustness of our numerical solver, we now consider fully two-way coupled FSI problems involving the sedimentation and buoyancy-driven motion of spherical particles in a quiescent fluid. These scenarios serve as a natural extension to the one-way coupled validation cases previously presented and pose additional numerical challenges due to the strong mutual interaction between the fluid and the solid dynamics. In our first numerical experiment, we simulate the free sedimentation of a spherical particle in a rectangular domain of size $10D\times 10D\times 15D$, where $D$ denotes the particle diameter. The particle is initially positioned along the vertical centerline at a height of $8.5D$ and released from rest. Free-slip boundary conditions were imposed on all faces of the computational domain to minimize wall effects and allow the particle to evolve as in an unbounded domain. Two grid resolutions, $200\times 200\times 300$ and $400\times 400\times 600$, were tested to assess sensitivity to spatial discretization. As shown in Fig. 9, the discrepancy between the solutions obtained on the coarse and fine grids does not exceed 7%, indicating satisfactory grid convergence. The simulated particle trajectories and velocities compare favorably with the experimental results reported by ten Cate et al. [56]. Since the fluid and solid dynamics are fully coupled, the simulations were carried out using the internal predictor-corrector iteration scheme described in Section 3.2.1, which is required to enforce the no-slip and divergence-free constraints at the fluid-solid interface. The iteration continues until the relative $\ell_{1}$-norm of the incremental force-density correction falls below a prescribed threshold, typically $\varepsilon=10^{-3}$. For all cases reported in this section, convergence was achieved within $2$ to $3$ iterations per time step. Importantly, the tested cases correspond to near-neutral buoyancy conditions, with density ratios $\rho_{p}/\rho_{f}\approx 1$, which are challenging due to the breakdown of the assumption that the fluid within the immersed body moves as a rigid body. In such regimes, the added-mass contributions must be explicitly captured for accurate prediction, further validating the correctness of our formulation.
(a) Time evolution of the vertical velocity
(b) Time evolution of the vertical coordinate
Figure 9: Sedimentation of a spherical particle within a quiescent fluid: present results for $\rho_{p}/\rho_{f}=1.164$, $Re_{p}=11.6$, $Fr=0.237$ and $\rho_{p}/\rho_{f}=1.167$, $Re_{p}=31.9$, $Fr=0.334$, compared against the experimental data of [56].
We further extend our validation by simulating additional two-way coupled cases involving both sedimenting and buoyant spherical particles. The computed vertical velocities are normalized using $u_{\mathrm{ref}}=\sqrt{gD_{p}}$ and $t_{\mathrm{ref}}=\sqrt{D_{p}/g}$ for velocity and time, respectively. The results are compared against both experimental data [57] and previous numerical simulations [40], as shown in Fig. 10. Excellent agreement is observed across entire range of density ratios and Reynolds numbers, further successfully verifying our two-way coupled immersed boundary solver.
(a) Present results (solid lines); experimental data from [57] (pluses); simulation results from [40] (circles). Case: $\rho_{p}/\rho_{f}=2.56$, $Re_{p}=367$, $Fr=1.8$.
(b) Present results (solid and dashed lines) compared to numerical results from [40]. Case: $Re_{p}=367$, $Fr=1.8$.
Figure 10: Time evolution of the vertical velocity of sedimenting and buoyant spherical particles in quiescent fluid: comparison with prior experimental and numerical data.
6.2 Preconditioner efficiency
The aim of this subsection is to validate the theoretical result of Theorem 5.1 numerically, and to demonstrate that the low iteration count needed for the solution of the preconditioned system is nearly constant, for different scenarios.
In the first experiment, we verify the spectral properties of the preconditioned system on a 2D framework consisting of a stationary cylinder placed at the center of a lid-driven cavity. In Table 2, we measure the spectrum’s support $\lambda_{min},\lambda_{max}]$ and show that $\text{spec}(L^{-1}S_{p})\subseteq[1,2]$, as predicted theoretically. Since the scalar approximation of $C$ is crucial for the efficiency of our method, we also measure the spectrum’s support for $L^{-1}\tilde{S}_{p}$, with $\tilde{S}_{p}$ from Eq. (33). In this case, most of the eigenvalues satisfy $\lambda\in[1,2]$, and although some eigenvalues slightly exceed 2, the maximal eigenvalue tends to 2 as the grid resolution increases. It can be explained by that fact that the quality of the scalar approximation $C\approx\frac{1}{2}I$ improves with the grid resolution, as shown in [30].
Spectrum’s support for the preconditioned system
Grid size (cells) $24\times 24$ $48\times 48$ $96\times 96$ $192\times 192$
In the second experiment we compare the GMRES iteration count required for the solution of the original system $\tilde{S}_{p}p^{\prime}={\bf b}$ and the preconditioned system $L^{-1}\tilde{S}_{p}p^{\prime}=L^{-1}{\bf b}$. Note that during a full solution of the problem, the right hand side $\bf b$} changes at any time step. Here, we solved the systems for a random right hand side to demonstrate the improvement of the conditioning in terms of iteration count. The results shown in Fig. 11 demonstrate that the preconditioned system can be solved in just a few iterations regardless of the grid size, while the original system requires hundreds of iterations for very small grids and thousands for larger grids. This scalability property enables performing the simulations on dense 3D grids, as shown in the next experiment.
Figure 11: GMRES iteration count for the solution of a system with a random right-hand-side for the corresponding matrices.
In the third experiment we verify that the scalability shown above extends to 3D experiments with moving bodies, by demonstrating it on model problem 1 from Subsection 6.1.1: an oscillating sphere. We evaluate the efficiency of the proposed preconditioner by measuring the average number of BiCGStab iterations required to converge the preconditioned system to a tolerance of $10^{-12}$ for various Reynolds numbers and grid sizes. Table 3 summarizes these results, averaged over three amplitude-to-diameter ratios (A/D=0.5, 1.0, 1.5A/D=0.5,\ 1.0,\ 1.5) and four representative locations of the sphere along its periodic trajectory: bottom dead center, top dead center, and the midpoints during upward and downward motion. The results demonstrate that the average iteration count remains nearly constant—typically between 4 and 5—across all tested Reynolds numbers, grid sizes, and sphere locations. This robustness highlights the effectiveness and scalability of the preconditioner.
Iteration count for a 3D oscillating sphere
Grid size $Re=10$ $Re=20$ $Re=50$ $Re=100$ $Re=200$ $Re=300$
$50\times 50\times 75$ 4.0 4.0 4.0 4.0 4.0 4.0
$100\times 100\times 150$ 4.25 4.0 4.0 4.17 4.17 4.0
$200\times 200\times 300$ 4.33 4.25 4.17 4.33 4.33 4.67
$400\times 400\times 600$ 4.17 4.0 4.08 4.08 4.33 4.25
Table 3: Average BiCGStab iteration count required for convergence of the preconditioned system to a $10^{-12}$ tolerance. Results are averaged over different $A/D$ ratios and over four representative locations of the sphere along its oscillation cycle.
To further assess the computational performance, we measured the wall-clock time per time step and memory consumption as functions of grid size. These experiments were conducted on a standard Linux server equipped with 64 GB DDR3 shared memory and two Intel Xeon 12C processors (48 threads total). The results, shown in Fig. 12, were averaged over the same set of $A/D$ ratios and six Reynolds numbers (Re=10, 20, 50, 100, 200, 300Re=10,\ 20,\ 50,\ 100,\ 200,\ 300). The measured wall-clock time scales approximately as $O(N^{1.1})$, where $N$ is the number of unknowns. This scaling aligns well with the expected complexity of the direct solver used in our implementation estimated as $O(N^{1.33})$ in [42, 43], which is further improved through the use of the Thomas algorithm in one coordinate direction. Memory consumption is sub-linear and remains modest: even for the largest grid ($400\times 400\times 600$, hosting approximately 384 million unknowns), the total memory requirement was only 25 GB. This is a significant improvement over previous work [47], where memory consumption exceeded 128 GB for comparable problem sizes at low Reynolds numbers. Importantly, the computational efficiency and memory consumption are nearly independent of the Reynolds number and time step, provided that the CFL number remains below 0.1.
(a) Time scalability
(b) Memory scalability
Figure 12: Wall-clock time per time step and memory consumption as functions of grid size. Results are averaged over three $A/D$ ratios (0.5, 1.0, 1.50.5,\ 1.0,\ 1.5) and six Reynolds numbers ($10$ to $300$). The solid line represents a linear regression trend.
@@ 6.2 Preconditioner efficiency
In the last experiment, we demonstrate the same scalability property on model problem 2 from Subsection 6.1.2, which contains multiple spheres. In the view of Remark 5.1, the number of immersed bodies might affect the sparsity of $B$, and thus the scattering of the eigenvalues of the preconditioned system. The aim of this experiment is to demonstrate that the same scalable behaviour holds, regardless of the number of bodies. Table 4 shows the BiCGStab iteration count needed to solve Eq. (31) with preconditioning, averaged over a single oscillation period, for different Reynolds numbers (Re=50, 100, 200, and 300) and grid resolutions. The results show that this period-averaged iteration count remains nearly constant, requiring 4-5 iterations for both 7 and 14 sub-sphere configurations. Furthermore, these period-averaged iteration counts match the results in Table 3 for a single oscillating sphere, demonstrating consistent performance across different configurations. These numerical results, combined with Theorem 5.1, demonstrate our method’s applicability and scalability across a wide range of applications.
@@ 7 Conclusion and future work
We present a novel, preconditioned direct-forcing IBM for efficient moving boundary and two-way coupled FSI simulations, building upon the SIMPLE-IBM framework [30]. To address the coupling between pressure and force-density corrections, we employ a block reduction technique that solves for the pressure correction first, using the Laplacian operator as a preconditioner for the resulting primal Schur complement. We rigorously prove and numerically verify that the Laplacian is spectrally equivalent to the primal Schur complement. This key observation enables an efficient and robust solution strategy: since the pressure Laplacian is independent of the immersed body configuration, it can be factorized once using a fast direct solver [42] and reused throughout the simulation.
@@
We demonstrate both theoretically and numerically that the performance of our solver is independent of grid resolution and the number of moving bodies, ensuring excellent scalability in terms of computational time and memory consumption. As a result, the method maintains low memory requirements, enabling high-fidelity moving boundary simulations that would typically require high-performance computing to run efficiently on standard workstations or even laptops. This improves the accessibility of advanced immersed boundary simulations for the broader CFD community. In this work, we demonstrated the applicability of the method to both moving boundary and two-way coupled FSI problems. The methodology was successfully verified by simulating flows induced by a transversely oscillating sphere, as well as sedimenting and buoyant spherical particles, showing good agreement with experimental and numerical reference data across a wide range of operating parameters. The approach was then extended to more complex configurations involving multiple moving bodies, specifically porous spheres modeled as arrays of rigid sub-spheres. The simulations revealed distinct flow features associated with different array porosities, including variations in drag coefficients, phase shifts, and vortex evolution, while preserving numerical stability and accuracy.
Future research should focus on extending the method to more complex two-way coupled FSI problems, particularly for simulating undulatory locomotion and deformable mesoscale porous media. The efficiency of our preconditioner does not depend on the specific choice of discrete delta function, which opens opportunities to improve accuracy by using wider-support kernels [10]. However, since our method relies on a scalar approximation of the regularization block, a key challenge will be to identify discrete delta functions with sufficiently wide support for which this scalar approximation remains valid.
Finally, the current implementation employs a direct Poisson solver [42] designed for homogeneous boundary conditions. While this is not restrictive in our context, where the pressure correction equation naturally adopts such conditions, future extensions may benefit from replacing the direct solver with a geometric multigrid V-cycle [58]. Such a strategy, if properly optimized, may achieve linear-time scaling and further enhance the versatility and accuracy of our approach for a wider class of FSI problems.